The ratio of number of electronic states in the energy range 2-3 eV to that of in 0-3 eV ((Number of electronic states in the energy range 2 -3 eV)/(Number of electronic states in the energy range 0-3 eV)) will be


1
Expert's answer
2022-02-25T11:12:57-0500

"p=\frac{n_{2-3}}{n_{0-3}}=1-\frac{2^2}{3^2}=0.56."